Publications and Preprints of J.P.C.Greenlees

[1] J.P.C.Greenlees “Functional duals and Moore spectra”, Bulletin of the London Mathematical Society 17 (1985), 43-48.

[2] J.P.C.Greenlees “Representing Tate cohomology of G-spaces”, Proceedings of the Edinburgh Mathematical Society 30 (1987), 435-443

[3] J.P.C.Greenlees “How blind is your favourite cohomology theory?”, Expositiones Mathematicae, 6 (1988), 193-208.

[4] J.P.C.Greenlees “Stable maps into free G-spaces”, Transactions of the American Mathematical Society, 310 (1988), 199-215.

[5] J.P.C.Greenlees “Topological methods in equivariant cohomology”, Proceedings of the 1987 Singapore Group Theory Conference, W. de Gruyter, (1989), 373-389.

[6] J.P.C.Greenlees “Equivariant functional duals and universal spaces”, Journal of the London Mathematical Society, 40 (1989), 347-354.

[7] J.P.C.Greenlees “The power of mod p Borel homology”, Proceedings of the 1988 Kinosaki conference on homotopy theory and related topics. Lecture notes in maths. 1418 Springer-Verlag (1990), 140-151.

[8] J.P.C.Greenlees “Equivariant functional duals and completions”, Bull. London Math. Soc. 23 (1991), 163-168.

[9] J.P.C.Greenlees and J.P. May “Completions of G-spectra at ideals of the Burnside ring” Proc. Adams Memorial Conference II, Cambridge University Press (1992) 145-178.

[10] J.P.C.Greenlees and J.P. May “Some remarks on the structure of Mackey functors” Proc. American Math. Soc., 395 (1992) 237-243.

[11] J.P.C.Greenlees and J.P. May “Derived functors of I-adic completion and local homology” Journal of Algebra 149 (1992) 438-453.

[12] J.P.C.Greenlees “Homotopy equivariance, strict equivariance and induction theory” Proc. Edinburgh Math. Soc. 35 (1992) 473-492.

[13] J.P.C.Greenlees “Generalised Eilenberg-Moore spectral sequences for elementary abelian groups and tori”, Proc. Camb. Phil. Soc., 112 (1992) 77-89.

[14] J.P.C.Greenlees “Some remarks on projective Mackey functors”, J. Pure and Applied Algebra, 81 (1992) 17-38.

[15] D.J. Benson and J.P.C.Greenlees “The action of the Steenrod algebra on Tate cohomology” J. Pure and Applied Algebra, 85 (1993) 21-26.

[16] J.P.C.Greenlees “K-homology of universal spaces and local cohomology of the representation ring” Topology 32 (1993) 295-308.

[17] T. Bier and J.P.C.Greenlees “The lattice spanned by the cosets of subgroups in a finite group.” J. London Math. Soc. 47 (1993) 433-449.

[18] J.P.C.Greenlees “The geometric equivariant Segal conjecture for toral groups.” J. London Math. Soc., 48 (1993) 348-364.

[19] A.D.Elmendorf, J.P.C.Greenlees, I.Kriz and J.P.May “Commutative algebra in stable homotopy theory and a completion theorem.” Mathematical Research Letters 1 (1994) 225-239.

[20] J.P.C.Greenlees “Tate cohomology in commutative algebra.” J. Pure and Applied Algebra 94 (1994) 59-83

[21] J.P.C.Greenlees and J.P. May “Generalized Tate cohomology” Memoirs of the American Maths. Soc., 543 (1995) 178pp.

[22] J.P.C.Greenlees “Commutative algebra in group cohomology.” J.Pure and Applied Algebra 98 (1995) 151-162

[23] J.P.C.Greenlees and J.P.May “Completions in algebra and topology” Handbook of Topology (ed. I.M.James) North Holland (1995) 255-276.

[24] J.P.C.Greenlees and J.P.May “Equivariant stable homotopy theory.” Handbook of Topology (ed. I.M.James) North Holland (1995) 277-323.

[25] R.Bruner and J.P.C.Greenlees “The algebraic Bredon-Löffler conjecture.” Experimental Mathematics 4 (1995) 289-297.

[26] J.P.C.Greenlees “A rational splitting theorem for the universal space for almost free actions.” Bull. London Math. Soc. 28 (1996) 183-189.

[27] J.P.C.Greenlees and H. Sadofsky “The Tate spectrum of v_{n}-periodic complex oriented theories.” Math. Zeits. 222 (1996) 391-405.

[28] J.P.C.Greenlees “An introduction to equivariant K-theory.” CBMS Regional Conference Series 91 American Math. Soc. (1996) 143-152.

[29] J.P.C.Greenlees and J.P.May “Examples of Tate cohomology.” CBMS Regional Conference Series 91 American Math. Soc. (1996) 231-245.

[30] J.P.C.Greenlees and J.P.May “Brave new equivariant algebra.” CBMS Regional Conference Series 91 American Math. Soc. (1996) 299-314.

[31] J.P.C.Greenlees and J.P.May “Localization and completion in complex bordism.” CBMS Regional Conference Series 91 American Math. Soc. (1996) 315-326.

[32] J.P.C.Greenlees and J.A.Pérez “Connected Lie groups that act freely on a product of spheres.” Bull. London Math. Soc. 28 (1996) 634-642.

[33] D.J.Benson and J.P.C.Greenlees “Commutative algebra for cohomology rings of classifying spaces of virtual duality groups.” J.Algebra 192 (1997) 678-700.

[34] D.J.Benson and J.P.C.Greenlees “Commutative algebra for cohomology rings of classifying spaces of compact Lie groups.” J. Pure and Applied Algebra 122 (1997) 41-53.

[35] J.P.C.Greenlees and H.Sadofsky “Tate cohomology of theories with one-dimensional coefficient ring.” Topology 37 (1998) 279-292.

[36] J.P.C.Greenlees and J.P.May “Localization and completion theorems for MU-module spectra.” Annals of Maths. 146 (1997) 509-544

[37] J.P.C.Greenlees “Rational Mackey functors for compact Lie groups I” Proc. London Math. Soc 76 (1998) 549-578

[38] J.P.C.Greenlees “Augmentation ideals of equivariant cohomology rings.” Topology 37 (1998) 1313-1323 [39] J.P.C.Greenlees “Rational O(2)-equivariant cohomology theories.” Fields Institute Communications

19 (1998) 103-110

[40] J.P.C.Greenlees “Rational S^{1}-equivariant stable homotopy theory.” Mem. American Math. Soc. 661 (Vol 138) (1999) vii+289 pp.

[41] J.P.C.Greenlees “Equivariant forms of connective K-theory.” Topology 38 (1999) 1075-1092.

[42] J.P.C.Greenlees and N.P.Strickland “Varieties and local cohomology for chromatic group cohomology rings.” Topology 38 (1999) 1093-1139.

[43] J.P.C.Greenlees and G.Lyubeznik “Rings with a local cohomology theorem and applications to cohomology rings of groups.” J. Pure and Applied Algebra 149 (2000) 267-285.

[44] M.M.Cole, J.P.C.Greenlees and I. Kriz “Equivariant formal group laws.” Proc. London Math. Soc 81 (2000) 355-386.

Accepted for publication:

[45] M.M.Cole, J.P.C.Greenlees and I. Kriz “Universality of equivariant bordism.” Math. Z. (to appear) 16pp

[46] J.P.C.Greenlees “Multiplicative equivariant formal group laws.” J. Pure and Applied Algebra (to appear) 20pp.

[47] J.P.C.Greenlees “Rational SO(3)-equivariant cohomology theories.” Proc Boulder Proceedings (to appear) 19pp

ON THE WAY TO PUBLICATION

Submitted for publication:

[1s] J.P.C.Greenlees “Tate cohomology in axiomatic stable homotopy theory” Preprint 16pp.

[2s] J.P.C.Greenlees, M.J.Hopkins and I. Rosu “Rational S^{1}-equivariant elliptic cohomology.” Preprint (1999) 23pp.

[3s] W.G.Dwyer and J.P.C.Greenlees “The equivalence of categories of complete and torsion modules.” Preprint (1999) 20pp

[4s] J.P.C.Greenlees “Local cohomology in equivariant topology.” Preprint (2000) 30pp

Preprints:

[1p] J.P.C.Greenlees “Equivariant formal groups over tame rings.” Preprint (1999) 15pp.

[2p] R.R.Bruner and J.P.C.Greenlees “The connective K-theory of finite groups.” Preprint (1998) 124pp.

[3p] J.P.C.Greenlees “Rational torus equivariant cohomology theories I: the standard model.” Preprint (2000) (20pp)

In preparation:

[4p] J.P.C.Greenlees “The coefficient ring of equivariant homotopical bordism is the universal ring for equivariant formal group laws.” (1998) 18pp

[5p] J.P.C.Greenlees “Rational Mackey functors for compact Lie groups II” 17pp.

[6p] J.P.C.Greenlees “Rational torus equivariant cohomology theories II: Adams spectral sequences.” (10pp)

[7p] J.P.C.Greenlees and B.E.Shipley “Rational torus equivariant cohomology theories III: Quillen equivalence with the standard model.” Preprint (1999) 8pp

[8p] J.P.C.Greenlees “Equivariant connective K theory for compact Lie groups” (In preparation) 10pp [9p] W.G.Dwyer, J.P.C.Greenlees and S.B.Iyengar “Gorenstein conditions in algebra and topology.” (In

preparation) 14pp

Inactive preprints:

[1i] J.P.C.Greenlees “Groups and Spheres: a road from the PA Smith theorem” (1992) 13pp.

[2i] J.P.C.Greenlees “A remark on local cohomology for non-Noetherian rings” (1992) 5pp.

THESIS

[1t.] J.P.C.Greenlees “Adams Spectral Sequences in equivariant topology”, Thesis, Cambridge University (1985), 381pp.

JPCG

September 2000